### v1

# Large Language Model Can Be a Foundation for Hidden Rationale-Based Retrieval

Abstract
Despite the recent advancement in Retrieval-Augmented Generation (RAG) systems, most retrieval methodologies are often developed for factual retrieval, which assumes query and positive documents are semantically similar. In this paper, we instead propose and study a more challenging type of retrieval task, called hidden rationale retrieval, in which query and document are not similar but can be inferred by reasoning chains, logic relationships, or empirical experiences. To address such problems, an instruction-tuned Large language model (LLM) with a cross-encoder architecture could be a reasonable choice. To further strengthen pioneering LLM-based retrievers, we design a special instruction that transforms the retrieval task into a generative task by prompting LLM to answer a binary-choice question. The model can be fine-tuned with direct preference optimization (DPO). The framework is also optimized for computational efficiency with no performance degradation. We name this retrieval framework by RaHoRe and verify its zero-shot and fine-tuned performance superiority on Emotional Support Conversation (ESC), compared with previous retrieval works. Our study suggests the potential to employ LLM as a foundation for a wider scope of retrieval tasks. Our codes, models, and datasets are available on https://github.com/flyfree5/LaHoRe.
Keywords:
LLM Retrieval RAG Hidden Rationale Generative Learning.1 Introduction
Retrieval-Augmented Generation (RAG) has demonstrated remarkable progress based on the development of the Large Language Model (LLM), which combines textual retrieval and data-augmented generation, to improve the response’s professionalism, controllability, explainability, and reduce hallucination [5, 28, 4, 29]. Given the user query, a retriever is obliged to select the most relevant document from the document corpus, to prepend into the generated prompt. Relevance between query and document can usually be drawn from semantic textual similarity (STS), in which the frequency-based retrieval methods, or bi-encoder architecture of dense retrievers can be applied. Such retrieval methodologies have been proven to be quite effective in traditional factual or knowledge-based question-answering systems (Fig. 1 (Left)).
However, wide applications of RAG systems sometimes challenge LLM for higher levels of cognitive capabilities, in which query and document are semantically related but not paraphrases of each other [23], and retrieval needs to be inferred from hidden rationale instead of explicitly factual-based [29, 25]. For example, one can develop a strategy pool and build a user-centric dialogue system that retrieves an appropriate response strategy given the user query [30, 3, 16]; a robot can search for the most valuable sub-task from plausible candidates determined by LLM, grounded by a global task instruction [1, 6, 19] (Fig. 1 (Right)). Although such problems can be viewed as or implemented with retrieval-based systems, traditional similarity-based retrievers might fail, since the relevance reasoning now is built upon underlying logical congruence or thematic alignment. Instead, LLMs with inner-embedded and in-context knowledge might be crucial with external data insufficiency.
Recent efforts of LLM-based retrieval either keep the bi-encoder architecture [11, 24, 23] (Fig. 2 (A)), or processes the query-document concatenation by cross-encoding [12, 8] (Fig. 2 (B)), or even combined with generative loss or task [13, 27] (Fig. 2 (C)). However, most of these efforts are still embedding-based approaches with the representing learning component achieved by contrastive loss from in-batch negatives or hard negatives [9]. Inspired by recent successful applications of generation tasks on retrieval studies, it is argued that an LLM-based cross-encoder with a solely generative loss might be able to understand the non-explicit semantic relation between query and document [14], becoming a foundation for hidden rationale-based retrieval problems.
In this paper, we propose a novel approach called LLM-based Hidden Rationale-based Retriever (LaHoRe), to address the aforementioned challenges. We combine the ideas of cross-encoder and the generative architectures, to conduct a paradigm shift from discriminative to generative modeling (Fig. 2 (D)). An instruction is appended to prompt a binary-choice question on the retrieval relevance, while the relevance score can be extracted from the choices’ relative log probability. To validate the effectiveness of LaHoRe, we construct several datasets from the field of Emotional Support Conversation (ESC) [10] and verify that LaHoRe outperforms previous baselines on the supporting strategy retrieval. The zero-shot performance of LaHoRe can be further improved by SFT or DPO. The main contributions include:
We expand the scope of traditional retrieval with a more challenging category called hidden rationale retrieval, and experiment it in emotional support conversations.
We propose a pure generative-based cross-encoder framework and convert the conventional contrastive loss into preference learning, with state-of-the-art performance obtained.
We intentionally switch the order of query and document and design the instruction correspondingly to obtain a computationally efficient retriever.
In the following contexts, we first recap the preliminaries of the semantic retrieval tasks, and then introduce the method details. After that, we illustrate the experiment results, including the datasets, zero-shot performance, fine-tuned performance, end-to-end RAG performance, ablation and sensitivity studies. We then discuss the related works and finally conclude the paper.
2 Method
In this section, we first introduce our prompt format, then propose the inference logic that provides the zero-shot result. Finally, the fine-tuning approach is discussed which further improves the retrieval accuracy. Fig. 2 (D) visualizes our framework.
2.1 Prompt Format
Given a query and a document , we design a special instruction, named , appended to the end of query-document concatenation:
The purpose of instruction is to transform the retrieval task into a generation task, which contains two parts, :
= ”Can Q be appropriately responded with D?”
= ”If you think the answer is true, please choose ; otherwise choose .”
where connects and from the sense of semantic relationship, and encourages LLM to thoroughly consider its preference on retrieval relevance by augmenting a binary-choice problem [26], and finally provide the answer, denoted by . An instruction fine-tuned LLM can constrain its by (””, ””), as expected.
(A): The bi-encoder architecture;
(B): The cross-encoder architecture;
(C): The generative RAG framework;
(D): Framework of LaHoRe. Q and D are switched with D’s decoding can be cached.
2.2 Inference
The relevance score can be obtained from the positive-negative relative confidence:
where denotes the next-token log probability from LLM. Although simple, we find this pipeline performs well even on zero-shot accuracy. Nevertheless, it is also computationally expensive since it requires a full forward pass for each pair.
Conversion into a efficient retriever: To alleviate the computational complexity, we intentionally put before in Equation 1, namely , instead of the more semantically natural order (). Given the causal decoding mechanism of LLM 111the latter tokens’ decoding do not affect the former., the prefix decoding or cache-sharing technology can be utilized to calculate and store the key-value cache of each off-the-shelf, while only decoding the part of during real-time inference with the cache loaded:
| offline: | ||||
| online: | (3) |
2.3 Annotation and Fine-Tuning
To leverage the scaling capability of LLM, each query is combined with all mismatched documents from the pool, to create the negative pairs. In the experiment section, we call this positive-negative ratio by ‘1:all’. The labels of are then annotated with ”” for positive pairs and ”” for negative pairs.
3 Experiment
In this section, we first introduce the datasets we use, then list zero-shot and fine-tuning results. Ablation and sensitivity studies are also conducted.
| Dataset() | ChatIntent | ChatStrategy | ESconv | PsyQA | ||||||||||
| Method () | Size | R@5 | R@10 | p-MRR | R@5 | R@10 | p-MRR | R@1 | R@3 | p-MRR | R@1 | R@3 | p-MRR | |
| Bi-Encoder | BM25[20] | - | 5.4 | 11.2 | 5.6 | 6.2 | 11.4 | 5.7 | 7.3 | 30.4 | 31.6 | 33.3 | 55.2 | 48.2 |
| TART-Contriever[2] | 110M | 4.2 | 9.2 | 5.1 | 4.8 | 8.9 | 5.1 | 1.4 | 38.5 | 35.1 | 11.3 | 44.2 | 36.5 | |
| E5-large-instruct[23] | 550M | 28.2 | 40.0 | 21.1 | 11.0 | 18.9 | 9.4 | 16.4 | 45.4 | 39.0 | 16.3 | 48.1 | 39.9 | |
| InstructOR XL[21] | 1.5B | 2.9 | 6.8 | 5.2 | 5.4 | 10.4 | 6.9 | 13.7 | 44.3 | 36.4 | 24.6 | 61.2 | 48.1 | |
| RepLLaMA[11] | 7B | 14.9 | 23.5 | 11.8 | 6.2 | 11.6 | 6.2 | 11.5 | 40.2 | 33.9 | 26.8 | 47.1 | 45.6 | |
| Promptriever[24] | 7B | 22.2 | 32.7 | 17.3 | 7.0 | 13.3 | 6.5 | 12.0 | 42.7 | 35.4 | 25.3 | 49.1 | 45.0 | |
| E5-Mistral[23] | 7B | 27.6 | 39.5 | 19.9 | 9.8 | 18.3 | 9.5 | 11.5 | 40.5 | 34.6 | 12.5 | 45.6 | 37.2 | |
| Cross-Encoder | LlaMA2Vec[8] | 7B | 15.9 | 26.6 | 12.0 | 4.8 | 10.0 | 5.4 | 16.4 | 44.4 | 38.0 | 28.0 | 49.4 | 46.7 |
| GritLM[13] | 7B | 29.5 | 40.2 | 22.3 | 14.4 | 23.9 | 11.0 | 15.9 | 45.8 | 38.5 | 5.3 | 37.9 | 30.8 | |
| LaHoRe (ours) | 7B | 30.7 | 48.2 | 22.1 | 16.5 | 28.3 | 11.6 | 16.9 | 42.7 | 38.0 | 32.6 | 52.5 | 49.5 | |
3.1 Datasets
Experimental datasets are constructed from either proprietary or open sources:
Emotional Support Conversation (ESC): we utilize two public datasets, ESconv [10] and PsyQA [22], with the name of ‘support strategies’ annotated. ESconv is a multi-turn dataset with 1202 dialogues, 12746 utterances, and 8 strategies. PsyQA is a single-turn dataset with 4012 dialogues and 7 strategies.
Daily Chat: we collect a proprietary multi-turn daily chat dataset with 1855 dialogues and 7910 utterances. Each user utterance is marked with an intent (e.g., ‘user asks for help’ or ’user disagrees’) and a reply strategy (e.g., ‘I need to show empathy’ or ’I should ask user’s habit’) by human annotators. Two retrieval datasets called ChatIntent and DailyStrategy are then created with 98 intents and 82 strategies totally.
For all four datasets, each user utterance is prepended by the historical context (if any) to form the query, with all possible strategies or intents as the documents.
3.2 Results
Considering numbers of documents of each dataset, here we mainly examine Recall@5, Recall@10, p-MRR for ChatIntent and DailyStrategy, and Recall@1, Recall@3, p-MRR for ESconv and PsyQA. LaHoRe is initialized by Qwen2-7B-Instruct [17]. Inference is implemented with vLLM with the prefix cache [7]. Training is conducted on LlamaFactory [31], with learning rate of , window length of 2048 and epoch of 3. Batch size is set as 512 for SFT and 128 for DPO. Train and test splits are divided by 9:1.
| Dataset() | ChatIntent | ChatStrategy | ESconv | PsyQA | |||||||||||
| Method () | R@1 | R@5 | R@10 | p-MRR | R@1 | R@5 | R@10 | p-MRR | R@1 | R@3 | p-MRR | R@1 | R@3 | p-MRR | |
| Bi-Encoder | RepLLaMA | 8.2 | 22.7 | 37.3 | 17.2 | 4.1 | 20.0 | 29.3 | 13.4 | 17.4 | 49.1 | 39.7 | 29.6 | 56.0 | 49.6 |
| Promptriever | 8.4 | 23.0 | 35.2 | 18.0 | 4.1 | 23.2 | 31.8 | 14.3 | 17.6 | 48.9 | 39.5 | 29.6 | 61.7 | 50.9 | |
| E5-Mistral | 31.1 | 59.5 | 71.8 | 44.9 | 28.4 | 61.1 | 74.8 | 43.8 | 15.5 | 50.2 | 39.3 | 11.4 | 48.0 | 36.8 | |
| Cross-Encoder | LlaMA2Vec | 35.7 | 71.1 | 80.7 | 50.6 | 37.5 | 76.4 | 85.0 | 53.8 | 23.9 | 55.6 | 45.5 | 34.3 | 75.4 | 57.5 |
| GritLM | 28.2 | 60.7 | 71.8 | 42.8 | 29.8 | 66.4 | 82.5 | 46.8 | 19.2 | 50.2 | 41.7 | 7.7 | 32.1 | 29.8 | |
| LaHoRe-SFT | 64.1 | 87.3 | 93.6 | 73.4 | 61.4 | 88.9 | 95.0 | 75.1 | 34.4 | 67.1 | 54.8 | 35.6 | 77.6 | 58.2 | |
| LaHoRe-DPO | 77.3 | 93.6 | 95.9 | 83.8 | 73.9 | 94.1 | 97.5 | 82.4 | 38.0 | 69.5 | 57.6 | 35.8 | 76.6 | 58.4 | |
Zero-shot Evaluation: Table 1 compares zero-shot results between Bi-Encoder and Cross-Encoder architectures. LLM-based baselines do not perform well and are even worse than smaller retrievers in some metrics. Our LaHoRe performs the (or the second) best across all metrics, which indicates LaHoRe can utilize knowledge and reasoning of LLM effectively, especially for more challenged tasks like ChatIntent and ChatStrategy.
Fine-Tuning Evaluation: Table 2 exhibits the fine-tuned results of LaHoRe and several 7B baselines. LaHoRe apparently outperforms other baselines, with the DPO version performing even better than SFT. This observation illustrates that generative preference optimization can be another retrieval basis besides the popular contrastive learning, and might be a stronger aligner to hidden rationale-based retrieval.
End-to-End RAG: We finally implement a RAG pipeline with Qwen2-70B-Instruct as the generator and DailyStrategy as the test environment. The retrieved reply strategy by LaHoRe is appended into the prompt to enhance LLM generation. Response qualities, including aspects of coherence, fluency and safety, are blindly evaluated by human annotators, and LaHoRe has a 90% win-rate in the comparison with raw LLM generation.
| Dataset() | ChatIntent | ChatStrategy | ESconv | PsyQA | RT | ||||||||||
| R@5 | R@10 | p-MRR | R@5 | R@10 | p-MRR | R@1 | R@3 | p-MRR | R@1 | R@3 | p-MRR | (ms) | |||
| LaHoRe | 10.3 | 18.0 | 8.9 | 4.9 | 8.3 | 4.9 | 17.3 | 37.9 | 36.5 | 19.5 | 47.4 | 35.7 | - | ||
| ✓ | 14.3 | 24.7 | 11.0 | 4.7 | 7.6 | 4.8 | 17.2 | 37.0 | 36.5 | 24.9 | 49.5 | 44.5 | - | ||
| ✓ | 31.5 | 50.0 | 19.9 | 10.2 | 21.4 | 8.6 | 18.1 | 43.9 | 39.2 | 30.8 | 52.0 | 48.3 | 47.5 | ||
| ✓ | ✓ | 30.0 | 48.2 | 22.1 | 16.5 | 28.3 | 11.6 | 16.9 | 42.7 | 38.0 | 32.6 | 52.5 | 49.5 | 19.3 | |
3.3 Analysis
Ablation Study: We conduct ablation study on two core features, and . Table 3 (the last row representing current LaHoRe) help answer several important questions:
RQ1: Result of raw LLM (the first row) is worse than not only LaHoRe but also many baselines in Table 1. This result rules out the capability difference of the basis model.
RQ2: With and without exhibit a huge difference which indicates its effectiveness.
RQ3: Performance ablation on is relative balanced. However, response time (RT) can be improved apparently with where the prefix-cache mechanism applies.
Sensitivity Analysis: In conventional embedding-based retrieval, it is common to construct in-batch negatives or hard negatives and train with contrastive loss. While in our generation-based paradigm, both positive or negative pairs are valid, independent training samples, therefore the positive-negative sample ratio could be an arbitrary choice. Compared with the formal choice of 1:all, we also experiment with other ratios including 1:1, 1:3 and 1.5, with plots of metrics shown in Fig. 3. It is evident that 1:all performs the best, which validates the data-scaling benefit of LaHoRe, even with positive outnumbered by negative samples. Also, this improvement is more apparent when the total number of documents is larger (ChatIntent and ChatStrategy).
4 Related Works
Most LLM-based retrieval studies still keep the bi-encoder architecture. For example, RepLLama [11] employed one Llama structure to encode the query and another to encode the document, and fine-tuned the retrieval task. Promptriever [24] added the prompt mechanism on RepLLama and obtained improved performance on instructed retrieval tasks. E5 [23] further improved the performance by generating and applying diverse, multilingual synthetic datasets. Such types of model architectures are aimed at solving the semantic similarity problem, which is significantly different from the hidden rationale retrieval problems proposed by this paper.
Recent efforts of LLM-based retrieval also attempt to cross-encoding the query and document, such that the non-similarity semantic problem can be studied. The ranking indicator might be obtained from the log probability of the query-document concatenation, such as [12], or still from the encoded query and document vectors [8]. There are also studies that start to turn the retrieval task into a combination of representative and generative problems [13, 27, 14]. On the other hand, to better study the hidden rationale retrieval problem, we propose a pure-generative framework, and append a binary choice question to the end of query-document concatenation such that its answer’s log probability can be employed as the ranking score directly. We also observe that the DPO loss in such a generative framework can be an analogy of contrastive loss in the traditional representing framework, and generates state-of-the-art performance in the problems we study in this work.
5 Conclusion
In this paper, we propose an LLM-based retrieval approach called LaHoRe, to address the challenging hidden rationale retrieval problem. We study the emotional supporting conversation scenario where different reply strategies are retrieved based on the user query. We leverage the reasoning and semantic understanding ability of LLM on this problem with solely generative loss. Specific instruction is designed to prompt LLM to answer a binary-choice question, and prefix decoding technology can be applied to make the retriever computationally efficient. Experimental results verify that both zero-shot and fine-tuned performances of LaHoRe surpass all other baselines on such types of tasks.
6 Limitation
The framework proposed in this paper, LaHoRe, is designed to solely address a special type of retrieval problem, the hidden rationale retrieval. It is expected that LaHoRe can not perform as well as conventional retrievers for semantic similarity retrieval tasks.
Furthermore, LaHoRe is a dense retrieval and might still suffer from computational burden as most cross-encoding approaches do. This issue is somewhat alleviated by our designed document-query order and the vLLM prefix cached inference, but still expected to be less computationally efficient than sparse retrievers or bi-encoder structures. Further works may include the knowledge distillation which converts the LLM knowledge into sparse encoders.

### v2

@@
 Keywords:
-LLM Retrieval RAG Hidden Rationale Generative Learning.1 Introduction
+LLM Hidden Rationale Retrieval RAG Generative.1 Introduction
 Retrieval-Augmented Generation (RAG) has demonstrated remarkable progress based on the development of the Large Language Model (LLM), which combines textual retrieval and data-augmented generation, to improve the response’s professionalism, controllability, explainability, and reduce hallucination [5, 28, 4, 29].
 Given the user query, a retriever is obliged to select the most relevant document from the document corpus, to prepend into the generated prompt.
-Relevance between query and document can usually be drawn from semantic textual similarity (STS), in which the frequency-based retrieval methods, or bi-encoder architecture of dense retrievers can be applied.
+Relevance between query and document can usually be drawn from semantic textual similarity (STS), in which the frequency-based retrieval methods [20], or bi-encoder architecture of dense retrievers [2, 21, 23] can be applied.
 Such retrieval methodologies have been proven to be quite effective in traditional factual or knowledge-based question-answering systems (Fig.
@@
 1 (Right)).
-Although such problems can be viewed as or implemented with retrieval-based systems, traditional similarity-based retrievers might fail, since the relevance reasoning now is built upon underlying logical congruence or thematic alignment.
-Instead, LLMs with inner-embedded and in-context knowledge might be crucial with external data insufficiency.
+Although such problems can be viewed as or implemented with retrieval-based systems, traditional similarity-based retrievers might fail, since the relevance reasoning now needs to be built upon underlying logical congruence or thematic alignment.
+Instead, LLMs with inner-embedded and in-context commonsense knowledge might be crucial tools, especially with external data insufficiency.
 Recent efforts of LLM-based retrieval either keep the bi-encoder architecture [11, 24, 23] (Fig.
@@
 2 (C)).
-However, most of these efforts are still embedding-based approaches with the representing learning component achieved by contrastive loss from in-batch negatives or hard negatives [9].
+However, most of these efforts are still embedding-based approaches, with the representing learning component achieved by contrastive loss of in-batch negatives or hard negatives [9].
 Inspired by recent successful application s of generation tasks on retrieval studies, it is argued that an LLM-based cross-encoder with a solely generative loss might be able to understand the non-explicit semantic relation between query and document [14], becoming a foundation for hidden rationale-based retrieval problems.
@@
 To validate the effectiveness of LaHoRe, we construct several datasets from the field of Emotional Support Conversation (ESC) [10] and verify that LaHoRe outperforms previous baselines on the supporting strategy retrieval.
+We also utilize the prompt cache technology to improve the computational efficiency by changing the order of query and document.
 The zero-shot performance of LaHoRe can be further improved by SFT or DPO.
-The main contributions include:
-We expand the scope of traditional retrieval with a more challenging category called hidden rationale retrieval, and experiment it in emotional support conversations.
-We propose a pure generative-based cross-encoder framework and convert the conventional contrastive loss into preference learning, with state-of-the-art performance obtained.
-We intentionally switch the order of query and document and design the instruction correspondingly to obtain a computationally efficient retriever.
-In the following contexts, we first recap the preliminaries of the semantic retrieval tasks, and then introduce the method details.
-After that, we illustrate the experiment results, including the datasets, zero-shot performance, fine-tuned performance, end-to-end RAG performance, ablation and sensitivity studies.
-We then discuss the related works and finally conclude the paper.
+The main contributions include the following:
+We expand the scope of traditional retrieval with a more challenging category called hidden rationale-based retrieval, and experiment it in the scenario of emotional support conversations.
+We design a pure generative-based cross-encoder framework and convert the conventional contrastive loss into preference learning, with state-of-the-art performance obtained.
+We propose a special solution to obtain a computationally efficient retriever.
 2 Method
-In this section, we first introduce our prompt format, then propose the inference logic that provides the zero-shot result.
-Finally, the fine-tuning approach is discussed which further improves the retrieval accuracy.
+In this section, we first introduce our prompt format, then propose the inference logic that provides zero-shot results.
+Finally the fine-tuning approach is discussed which further improves the performance.
 Fig.
@@
 The purpose of instruction is to transform the retrieval task into a generation task, which contains two parts, :
-= ”Can Q be appropriately responded with D?”
-= ”If you think the answer is true, please choose ; otherwise choose .”
-where connects and from the sense of semantic relationship, and encourages LLM to thoroughly consider its preference on retrieval relevance by augmenting a binary-choice problem [26], and finally provide the answer, denoted by .
-An instruction fine-tuned LLM can constrain its by (””, ””), as expected.
-(A): The bi-encoder architecture;
-(B): The cross-encoder architecture;
-(C): The generative RAG framework;
-(D): Framework of LaHoRe. Q and D are switched with D’s decoding can be cached.
+= "Can Q be appropriately responded with D?"
+= "If you think the answer is true, choose <T>; otherwise choose <F>."
+where connects and from the sense of semantic relation, and encourages LLM to thoroughly consider its preference on the relevance by augmenting a binary-choice problem [26], and finally provide the answer .
+An instruction fine-tuned LLM can constrain its by ("<T>", ""), as expected.
+(A): Bi-encoder; (B): Cross-encoder; (C): Generative RAG framework; (D): Framework of LaHoRe. We switch the order of Q and D such that D’s encoding can be cached.
 2.2 Inference
-The relevance score can be obtained from the positive-negative relative confidence:
-where denotes the next-token log probability from LLM.
-Although simple, we find this pipeline performs well even on zero-shot accuracy.
-Nevertheless, it is also computationally expensive since it requires a full forward pass for each pair.
-Conversion into a efficient retriever: To alleviate the computational complexity, we intentionally put before in Equation 1, namely , instead of the more semantically natural order ().
-Given the causal decoding mechanism of LLM 111the latter tokens’ decoding do not affect the former., the prefix decoding or cache-sharing technology can be utilized to calculate and store the key-value cache of each off-the-shelf, while only decoding the part of during real-time inference with the cache loaded:
-| offline: | ||||
-| online: | (3) |
+The relevance score is deduced from the positive-negative relative confidence:
+where denotes LLM’s next-token log probability.
+Although simple, we find this pipeline performs well even on zero-shot tests.
+However, it is computationally expensive since it requires a full forward pass for each pair.
+Conversion into a efficient retriever: To alleviate computational complexity, we intentionally put before in Equation 1, namely , instead of the more semantically natural order ().
+Given the causal decoding mechanism of LLM, the prefix decoding or cache-sharing technology [7] can be utilized to calculate and store the key-value cache of each off-the-shelf, while only decoding the part of during real-time inference with the cache loaded:
 2.3 Annotation and Fine-Tuning
@@
 In the experiment section, we call this positive-negative ratio by ‘1:all’.
-The labels of are then annotated with ”” for positive pairs and ”” for negative pairs.
+The labels of are then annotated with "<T>" for positive pairs and "<F>" for negative pairs.
 3 Experiment
 In this section, we first introduce the datasets we use, then list zero-shot and fine-tuning results. Ablation and sensitivity studies are also conducted.
+| All Strategies | [Question,Restatement or Paraphrasing,Reflection of feelings,Information, Self-disclosure,Affirmation and Reassurance,Providing Suggestions,Others] |
+| Query | {history} user: Seriously! What I am scare of now is how to secure another job. |
+| Strategy | Reflection of feelings |
+| Response | assistant: I can feel your pain just by chatting with you. |
+3.1 Datasets
+Experimental datasets are constructed from either proprietary or open sources:
+Emotional Support Conversation (ESC): we utilize two public datasets, ESconv [10] and PsyQA [22], with ‘support strategy’ annotated for each assistant utterance.
+ESconv is multi-turn with 1202 dialogues, 12746 utterances, and 8 strategies.
+PsyQA is single-turn with 4012 dialogues and 7 strategies.
+Daily Chat: we collect a proprietary multi-turn daily chat dataset including 1855 dialogues and 7910 utterances.
+Each turn is annotated with a user-side intent (e.g., ‘user asks for help’ or ’user disagrees’) and an assistant-side reply strategy (e.g., ‘I need to show empathy’ or ’I should ask user’s habit’) by human labelers.
+Two retrieval datasets called ChatIntent and DailyStrategy are then created with 98 intents and 82 strategies totally.
+For all four datasets, each user utterance is prepended by the historical context (if any) to form the query, with all possible strategies or intents as documents.
+One can refer to Table 1 for an exampled sample from ESconv.
+PsyQA and DailyStrategy have similar formats while ChatIntent has the intents as documents instead of strategies.
 | Dataset() | ChatIntent | ChatStrategy | ESconv | PsyQA | ||||||||||
@@
 | LaHoRe (ours) | 7B | 30.7 | 48.2 | 22.1 | 16.5 | 28.3 | 11.6 | 16.9 | 42.7 | 38.0 | 32.6 | 52.5 | 49.5 | |
-3.1 Datasets
-Experimental datasets are constructed from either proprietary or open sources:
-Emotional Support Conversation (ESC): we utilize two public datasets, ESconv [10] and PsyQA [22], with the name of ‘support strategies’ annotated.
-ESconv is a multi-turn dataset with 1202 dialogues, 12746 utterances, and 8 strategies.
-PsyQA is a single-turn dataset with 4012 dialogues and 7 strategies.
-Daily Chat: we collect a proprietary multi-turn daily chat dataset with 1855 dialogues and 7910 utterances.
-Each user utterance is marked with an intent (e.g., ‘user asks for help’ or ’user disagrees’) and a reply strategy (e.g., ‘I need to show empathy’ or ’I should ask user’s habit’) by human annotators.
-Two retrieval datasets called ChatIntent and DailyStrategy are then created with 98 intents and 82 strategies totally.
-For all four datasets, each user utterance is prepended by the historical context (if any) to form the query, with all possible strategies or intents as the documents.
 3.2 Results
-Considering numbers of documents of each dataset, here we mainly examine Recall@5, Recall@10, p-MRR for ChatIntent and DailyStrategy, and Recall@1, Recall@3, p-MRR for ESconv and PsyQA.
-LaHoRe is initialized by Qwen2-7B-Instruct [17].
-Inference is implemented with vLLM with the prefix cache [7].
-Training is conducted on LlamaFactory [31], with learning rate of , window length of 2048 and epoch of 3.
+The train-test split of datasets is 9:1.
+Considering the amount of documents, here we examine Recall@5, Recall@10, p-MRR for ChatIntent and DailyStrategy, and Recall@1, Recall@3, p-MRR for ESconv and PsyQA.
+LaHoRe is initialized from Qwen2-7B-Instruct [17].
+Training is conducted on LlamaFactory [31], with the learning rate of , window length of 2048 and epoch of 3.
 Batch size is set as 512 for SFT and 128 for DPO.
-Train and test splits are divided by 9:1.
 | Dataset() | ChatIntent | ChatStrategy | ESconv | PsyQA | |||||||||||
@@
 3.3 Analysis
-Ablation Study: We conduct ablation study on two core features, and . Table 3 (the last row representing current LaHoRe) help answer several important questions:
+Ablation Study: We conduct ablation study on two core features, and . Table 4 help answer several important questions:
 RQ1: Result of raw LLM (the first row) is worse than not only LaHoRe but also many baselines in Table 2. This result rules out the capability difference of the basis model.
@@
 Also, this improvement is more apparent when the total number of documents is larger (ChatIntent and ChatStrategy).
-4 Related Works
-Most LLM-based retrieval studies still keep the bi-encoder architecture.
-For example, RepLLama [11] employed one Llama structure to encode the query and another to encode the document, and fine-tuned the retrieval task.
-Promptriever [24] added the prompt mechanism on RepLLama and obtained improved performance on instructed retrieval tasks.
-E5 [23] further improved the performance by generating and applying diverse, multilingual synthetic datasets.
-Such types of model architectures are aimed at solving the semantic similarity problem, which is significantly different from the hidden rationale retrieval problems proposed by this paper.
-Recent efforts of LLM-based retrieval also attempt to cross-encoding the query and document, such that the non-similarity semantic problem can be studied.
-The ranking indicator might be obtained from the log probability of the query-document concatenation, such as [12], or still from the encoded query and document vectors [8].
-There are also studies that start to turn the retrieval task into a combination of representative and generative problems [13, 27, 14].
-On the other hand, to better study the hidden rationale retrieval problem, we propose a pure-generative framework, and append a binary choice question to the end of query-document concatenation such that its answer’s log probability can be employed as the ranking score directly.
-We also observe that the DPO loss in such a generative framework can be an analogy of contrastive loss in the traditional representing framework, and generates state-of-the-art performance in the problems we study in this work.
 4 Conclusion
@@
 Experimental results verify that both zero-shot and fine-tuned performances of LaHoRe surpass all other baselines on such types of tasks.
-6 Limitation
-The framework proposed in this paper, LaHoRe, is designed to solely address a special type of retrieval problem, the hidden rationale retrieval.
-It is expected that LaHoRe can not perform as well as conventional retrievers for semantic similarity retrieval tasks.
-Furthermore, LaHoRe is a dense retrieval and might still suffer from computational burden as most cross-encoding approaches do.
-This issue is somewhat alleviated by our designed document-query order and the vLLM prefix cached inference, but still expected to be less computationally efficient than sparse retrievers or bi-encoder structures.
-Further works may include the knowledge distillation which converts the LLM knowledge into sparse encoders.